# Towards Quantifying Long-Range Interactions in Graph Machine Learning: a Large Graph Dataset and a Measurement

### v2

@@ Abstract
Long-range dependencies are critical for effective graph representation learning, yet most existing datasets focus on small graphs tailored to inductive tasks, offering limited insight into long-range interactions. Current evaluations primarily compare models employing global attention (e.g., graph transformers) with those using local neighborhood aggregation (e.g., message-passing neural networks) without a direct measurement of long-range dependency. In this work, we introduce City-Networks, a novel large-scale transductive learning dataset derived from real-world city road networks. This dataset features graphs with over $10^{5}$ nodes and significantly larger diameters than those in existing benchmarks, naturally embodying long-range information. We annotate the graphs based on local node eccentricities, ensuring that the classification task inherently requires information from distant nodes. Furthermore, we propose a model-agnostic measurement based on the Jacobians of neighbors from distant hops, offering a principled quantification of long-range dependencies. Finally, we provide theoretical justifications for both our dataset design and the proposed measurement—particularly by focusing on over-smoothing and influence score dilution—which establishes a robust foundation for further exploration of long-range interactions in graph neural networks. Our dataset and measurement are now both available on PyTorch Geometric, and the code for reproducing the experimental results can be found at github.com/LeonResearch/City-Networks.
@@ 1 Introduction
Graphs are a widely used mathematical abstraction across nearly every branch of science. They are particularly effective for modeling the intricate and non-uniform interactions found in real-world data, where nodes stand in for objects and edges depict their connections. The growing recognition of the versatility of graph representations has sparked intense interest in Graph Neural Networks (GNNs) (Scarselli et al., 2009; Wu et al., 2021), driving innovation in deep learning for both geometric and graph-centric applications. Most GNNs, in particular variants based on a message-passing mechanism (Gilmer et al., 2017), exchange information between one-hop neighbors per layer to build node representations. While they found wide success in analysing citation and social networks (Yang et al., 2016b), one significant limitation concerns their capability of handling long-range dependencies, where interactions between distant nodes might be required to solve a task. Most existing datasets are not sufficient in assessing this: for instance, social networks, despite comprising thousands of nodes, often exhibit the small world effect (Watts & Strogatz, 1998) with short average path lengths and high clustering coefficients. The node labels on these graphs usually possess a homophilic pattern where nodes tend to connect with “similar” or “alike” others (McPherson et al., 2001), making it possible to propagate sufficient information for modelling with only a few message-passing layers. On the other hand, while connected nodes tend to have different properties on heterophilic graphs (Zhu et al., 2020; Ma et al., 2022), solving the tasks in those cases do not necessarily require the handling of long-range dependencies (Arnaiz-Rodriguez & Errica, 2025).
@@
Recently, Long Range Graph Benchmark (LRGB(Dwivedi et al., 2023) introduces alternative graph datasets based on super-pixels and molecules. To justify the existence of long-rangeness, the authors compare classical GNNs with Graph Transformers (GTs), which leverage global attentions over the entire graph, and associate the observed performance gaps with the presence of long-range dependencies. However, conclusions that are solely derived from empirical comparisons may not be reliable as they can be largely influenced by hyperparameter tuning (Tönshoff et al., 2024), leading to an ambiguous assessment of the long-range interactions. Moreover, LRGB and other synthetic long-range benchmarks (Bodnar et al., 2021; Zhou et al., 2025) all focus on inductive learning tasks with relatively small graphs—typically on the order of $10$ to $10^{2}$ nodes, while currently there is no long-range benchmark that considers large graphs with real-world topology for transductive learning. This is a critical gap in the literature since applying GTs (Rampášek et al., 2022), which are expected to better capture long-range interactions, to large graphs is significantly more challenging compared to small-graph inductive tasks due to the computational complexity of global attention (Borde, 2024).
@@
  • To quantify such long-range dependency, we further introduce a model-agnostic measurement that quantifies per-hop influence of a focal node’s neighbors on its prediction, which is estimated by the aggregated $\ell_{1}$-norm of the Jacobian from a trained model at each hop around the focal node. As a validation, we observe that distant hops exert a greater influence on all baseline models in our City-Networks compared to those on other commonly used graph datasets in the literature.
@@ 2.1 Large-scale Road Networks with Long Diameters
Real-world network topologies and features. As shown in Figure 1, our City-Networks consists of street maps in four cities: Paris, Shanghai, Los Angeles, and London, which are obtained by querying OpenStreetMap (Haklay & Weber, 2008) with OSMnx (Boeing, 2024). In particular, we consider a city network inclusive of all types of roads in the city (e.g., drive, bike, walk, etc.), where nodes represent road junctions with features like longitude, latitude, land use, etc.; and edges represent road segments with features like road length, speed limit, road types, etc. Next, to facilitate a typical node classification task, we apply a simple neighborhood aggregation that transforms edge features into node features by averaging the features of incidental edges and then concatenating them to the features of the focal node. These new node features represent, for instance, the average speed limit around a road junction or the probability of finding a residential road nearby. As a result, the final dataset contains $37$ node features, where we also make the graphs undirected so that they only represent connectivity between junctions, and retain the largest connected component in each city. We refer readers to Appendix B.1 for a more detailed discussion on our dataset processing approach.
@@ 2.2 Long-range labels based on urban accessibility
However, using the exact eccentricity as signals is not ideal for two reasons. First, we cannot control the level of long-rangeness, as eccentricity is determined by the entire graph structure and not adjustable by design. In particular, although predicting $\varepsilon(v)$ appears to require distant information based on its definition, in many cases (Figure 6, Appendix B.2), the underlying signal can be well inferred by spatial features (e.g., geographical coordinates) alone. This is because nodes in the city center generally have lower eccentricity due to the grid-like topology in road networks. Second, the calculation of eccentricity requires computing all pair-wise shortest paths in the network, which is typically at cost $\mathcal{O}(|V|^{3})$ and is hence prohibitively expensive for labeling our large-scale networks.
@@
We emphasize that this local approach is introduced not only for computational efficiency, but it also brings in two important considerations. First, it creates long-range signals that extend to $k$-hop neighbors by design, i.e., a known “ground-truth” range, hence is directly controllable and testable. Second, it ensures both graph topology and node features contribute to the modelling, avoiding the problem of high correlation between node signal with either the graph structure or node features alone. Indeed, since $k$ and the edge weights used to compute $\hat{\varepsilon}_{k}(v)$ are unknown to the model, it must integrate structural information from distant neighborhoods with their spatial information (geographic location, road type, land use, etc.) to infer the long-range signal, which makes the task non-trivial.
@@
Choice of the long-range level $k$. To assess long-range dependency, $k$ should be sufficiently large to distinguish our setting from short-range tasks (e.g., social graphs with typical message passing around $4$ hops). Thanks to the large grid-like topology with long diameters, we can create long enough node signals based on different local networks. At the same time, $k$ should also not be too large such that (i) the selected neighborhoods maintain local characteristics; and (ii) baseline models under a $k$-layer architecture, which captures the required information, can fit into common GPUs for fair benchmarking purposes. After experimenting with values from $8$ to $32$ (Figure 8, Appendix B.4), we design the task at kk=$16$, which strikes a balance between sufficient long-rangeness and computational efficiency. We refer readers to Appendix B.4 for a more detailed justification of our methods.
@@
Visualizations and interpretations. As shown in Equation (1), calculating $\hat{\varepsilon}_{k}(v)$ naturally requires information from distant neighbors up to $k$ hops, and it carries a practical meaning which relates to the accessibility of node $v$ in the network. As an illustration, we visualize two sub-regions from the city maps in Figure 2 with kk=$16$. We can observe that nodes on major transportation routes such as freeways and highways tend to have larger eccentricities than those in populated areas. This is because reaching a node’s 16-hop neighbors from a highway junction often requires traveling a much longer distance compared to road junctions in a downtown area. It is also clear from Figure 2 that a significant part of the graph topology is grid-like and possibly quasi-isometric to a lattice, as we will later discuss in Section 6. Note that the homophily scores, as reported in Table 1, are reasonably high since nodes with similar $\hat{\varepsilon}_{k}$ tend to cluster together.
@@
New challenges for learning long-range dependencies. The proposed dataset presents a significant challenge for both GNNs and GTs. In general, such long-range dependencies create a fundamental trade-off for graph learning: while increasing the number of layers in GNNs helps propagate information from distant hops, it also leads to issues such as over-smoothing (Li et al., 2018; Nt & Maehara, 2019; Rusch et al., 2023), over-squashing (Alon & Yahav, 2021; Topping et al., 2022; Di Giovanni et al., 2023), and vanishing gradients (Arroyo et al., 2025). On the other hand, GTs, which rely on attention mechanisms with quadratic computational complexity, face scalability challenges when applied to our large-scale city networks compared to learning on smaller long-ranged graphs like those in LRGB (Dwivedi et al., 2023) and other synthetic benchmarks (Zhou et al., 2025).
@@ 3 Benchmarks: From Classical GNNs to Graph Transformers
Experimental setups. We consider transductive node classification with train/validation/test splits of $10\%/10\%/80\%$ on all City-Networks, in which we evaluate various classical GNNs and GTs. Specifically, we benchmark GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCNII (Chen et al., 2020) for GNNs; and GraphGPS (Rampášek et al., 2022), Exphormer (Shirzad et al., 2023), and SGFormer (Wu et al., 2023) for GTs. In addition, an MLP is also included to reflect the importance of graph topology (or lack thereof) in our task. Lastly, we fix the data split for benchmarking purposes and report the mean and standard deviation of the test accuracy over 5 runs with different random seeds.
@@
For fair comparison purposes, we follow the latest GNN tuning technique from Luo et al. (2024) that considers residual connection, normalization, dropout, etc. Readers are also referred to Appendix C.2 for full details on our training and experimental setups.
@@
For completeness, we further conducted an ablation study in Appendix C.3 that (i) tests baselines under different hidden_size; and (ii) fixes the model size at $L=16$, refrains the model from seeing beyond hop-$H$ at each node, then tests baselines at different $H$. The results are all consistent with our above findings. In addition, we also summarize the best baseline performance on City-Networks in Appendix C.4, and discuss the potential limitation of our experiments in Appendix E.
@@ 4 A Long-range Measurement Based on Jacobians
Given Equation (2), we further define the average size $R$ of the influence-weighted receptive field as:
@@
where $H$ is the maximum number of hops to be considered. Intuitively, one can understand $R$ as measuring how far away the average unit of influence is, where graphs with long-range dependencies are expected to have higher proportional influence between more distant nodes. Note that while the proposed $R$ bears some similarity to the recent work (Bamberger et al., 2025), they do not consider per-hop influence $T_{h}(v)$ as a diagnostic tool of the dependency decay. Finally, we refer readers to Appendix D for a discussion on the computational cost of this measurement.
@@
To better compare $\bar{T}_{h}$ across models and datasets, we normalize it by $\bar{T}_{0}$ and report $\bar{T}_{h}/\bar{T}_{0}$ in Figure 4. For all models, we generally observe a rapid influence drop at distant hops on Cora and ogbn-arxiv, whereas the decay is noticeably much slower on City-Networks. The patterns on Amazon-Ratings and PascalVOC-SP generally fall between the above two cases with influence concentrated on the first few hops, which further corroborate our findings in Section 3. While $R$ and $\bar{T}_{h}$ are both model-dependent, the above results yield a consistent ordering across datasets: models trained on City-Networks are required to leverage information from more distant hops to perform well, compared to those trained on existing datasets. For the rest of the paper, we will first provide a theoretical justification of the proposed dataset by linking over-smoothing to algebraic connectivity and network diameter in Section 5, and then present the intuition behind the proposed long-range measurement in Section 6.

### v3

@@ Abstract
Long-range dependencies are critical for effective graph representation learning, yet most existing datasets focus on small graphs tailored to inductive tasks, offering limited insight into long-range interactions. Current evaluations primarily compare models employing global attention (e.g., graph transformers) with those using local neighborhood aggregation (e.g., message-passing neural networks) without a direct measurement of long-range dependency. In this work, we introduce City-Networks, a novel large-scale transductive learning dataset derived from real-world city road networks. This dataset features graphs with over $10^{5}$ nodes and significantly larger diameters than those in existing benchmarks, naturally embodying long-range information. We annotate the graphs based on local node eccentricities, ensuring that the classification task inherently requires information from distant nodes. Furthermore, we propose a generic measurement based on the Jacobians of neighbors from distant hops, offering a principled quantification of long-range dependencies. Finally, we provide theoretical justifications for both our dataset design and the proposed measurement—particularly by focusing on over-smoothing and influence score dilution, which establishes a robust foundation for further exploration of long-range interactions in graph neural networks. Our dataset and measurement are available on PyTorch Geometric, and the code for reproducing the experimental results can be found at github.com/LeonResearch/City-Networks.
@@ 1 Introduction
Graphs are a widely used mathematical abstraction across nearly every branch of science. They are particularly effective for modeling the intricate and non-uniform interactions found in real-world data, where nodes stand in for objects and edges depict their connections. The growing recognition of the versatility of graph representations has sparked intense interest in Graph Neural Networks (GNNs) (Scarselli et al., 2009; Wu et al., 2021), driving innovation in deep learning for both geometric and graph-centric applications. Most GNNs, in particular variants based on a message-passing mechanism (Gilmer et al., 2017), exchange information between one-hop neighbors per layer to build node representations. While they found wide success in analysing citation and social networks (Yang et al., 2016b), one significant limitation concerns their capability of handling long-range dependencies, where interactions between distant nodes might be required to solve a task. Most existing datasets are not sufficient in assessing this: for instance, social networks, despite comprising thousands of nodes, often exhibit the small world effect (Watts and Strogatz, 1998) with short average path lengths and high clustering coefficients. The node labels on these graphs usually possess a homophilic pattern where nodes tend to connect with “similar” or “alike” others (McPherson et al., 2001), making it possible to propagate sufficient information for modelling with only a few message-passing layers. On the other hand, while connected nodes tend to have different properties on heterophilic graphs (Zhu et al., 2020; Ma et al., 2022), solving the tasks in those cases do not necessarily require the handling of long-range dependencies (Arnaiz-Rodriguez and Errica, 2025).
@@
Recently, Long Range Graph Benchmark (LRGB(Dwivedi et al., 2023) introduces alternative graph datasets based on super-pixels and molecules with larger diameters than those of the previous works. To justify the existence of long-rangeness, the authors compare classical GNNs with Graph Transformers (GTs), which leverage global attentions over the entire graph, and associate the observed performance gaps with the presence of long-range dependencies. However, conclusions that are mainly derived from empirical comparisons may not be reliable as they can be largely influenced by hyperparameter tuning (Tönshoff et al., 2024), leading to an ambiguous assessment of the long-range interactions. Moreover, LRGB and other synthetic long-range benchmarks (Bodnar et al., 2021; Zhou et al., 2025) all focus on inductive learning tasks with relatively small graphs—typically on the order of $10$ to $10^{2}$ nodes, while currently there is no long-range benchmark that considers large graphs with real-world topology for transductive learning. This is a critical gap in the literature since applying GTs (Rampášek et al., 2022), which are expected to better capture long-range interactions, to large graphs is significantly more challenging compared to small-graph inductive tasks due to the computational complexity in its positional encoding and global attention (Borde, 2024).
@@
  • To quantify such long-range dependency, we further introduce a generic [^1] Here, generic means that the metric can be applied to any differentiable GNN or Graph Transformer, without requiring architectural modifications or assumptions about the model class. measurement that quantifies per-hop influence of a focal node’s neighbors on its prediction, which is estimated by the aggregated $\ell_{1}$-norm of the Jacobian from a trained model for the task at hand at each hop around the focal node. This per-hop analysis of the task range goes beyond the concurrent work of Bamberger et al. (2025) and provides novel insights: we observe that distant hops exert a greater influence on all baseline models in our City-Networks compared to those on other commonly used graph datasets in the literature.
@@ 2.1 Large-scale Road Networks with Long Diameters
Real-world network topologies and features. As shown in Figure 1, our City-Networks consists of street maps in four cities: Paris, Shanghai, Los Angeles, and London, which are obtained by querying OpenStreetMap (Haklay and Weber, 2008) with OSMnx (Boeing, 2024). In particular, we consider a city network inclusive of all types of roads in the city (e.g., drive, bike, walk, etc.), where nodes represent road junctions with features like longitude, latitude, land use, etc.; and edges represent road segments with features like road length, speed limit, road types, etc. Next, to facilitate a typical node classification task, we apply a simple neighborhood aggregation that transforms edge features into node features by averaging the features of incidental edges and then concatenating them to the features of the focal node. These new node features represent, for instance, the average speed limit around a road junction or the probability of finding a residential road nearby. As a result, the final dataset contains $37$ node features, where we also make the graphs undirected so that they only represent connectivity between junctions, and retain the largest connected component in each city. We refer readers to Appendix C.1 for a complete list of features and our approach in dataset processing.
@@ 2.2 Long-range labels based on urban accessibility
However, using the exact eccentricity as signals is not ideal for two reasons. First, we cannot control the level of long-rangeness, as eccentricity is determined by the entire graph structure and not adjustable by design. In particular, although predicting $\varepsilon(v)$ appears to require distant information based on its definition, in many cases (Figure 7, Appendix C.2), the underlying signal can be well inferred by spatial features (e.g., geographical coordinates) alone. This is because nodes in the city center generally have lower eccentricity due to the grid-like topology in road networks. Second, the calculation of eccentricity requires computing all pair-wise shortest paths in the network, which is typically at cost $\mathcal{O}(|V|^{3})$ and is hence prohibitively expensive for labeling our large-scale networks.
@@
We emphasize that this local approach is introduced not only for computational efficiency, but it also brings in two important considerations. First, it creates long-range signals that extend to $k$-hop neighbors by design, i.e., a known “ground-truth” range, hence is directly controllable and testable. Second, it ensures both graph topology and node features contribute to the modelling, avoiding the problem of high correlation between node signal with either the graph structure or node features alone. Indeed, since $k$ and the edge weights used to compute $\hat{\varepsilon}_{k}(v)$ are unknown to the model, it must integrate structural information from distant neighborhoods with their spatial information (geographic location, road type, land use, etc.) to infer the long-range signal, which makes the task non-trivial. We provide a deeper analysis in Appendix C.3 regarding the roles of graph structure and spatial features.
@@
Choice of the long-range level $k$. To assess long-range dependency, $k$ should be sufficiently large to distinguish our setting from short-range tasks (e.g., social graphs with typical message passing around $4$ hops). Thanks to the large grid-like topology with long diameters, we can create long enough node signals based on different local networks. At the same time, $k$ should also not be too large such that (i) the selected neighborhoods maintain local characteristics; and (ii) baseline models under a $k$-layer architecture, which captures the required information, can fit into common GPUs for fair benchmarking purposes. After experimenting with values from $8$ to $32$ (Figure 9, Appendix C.4), we design the task at kk=$16$, which strikes a balance between sufficient long-rangeness and computational efficiency. We refer readers to Appendix C.4 for a more detailed justification of our methods.
@@
Visualizations and interpretations. As shown in equation 1, calculating $\hat{\varepsilon}_{k}(v)$ naturally requires information from distant neighbors up to $k$ hops, and it carries a practical meaning which relates to the accessibility of node $v$ in the network. As an illustration, we visualize two sub-regions from the city maps in Figure 2 with kk=$16$. We can observe that nodes on major transportation routes such as freeways and highways tend to have larger eccentricities than those in populated areas. This is because reaching a node’s 16-hop neighbors from a highway junction often requires traveling a much longer distance compared to road junctions in a downtown area. It is also clear from Figure 2 that a significant part of the graph topology is grid-like and possibly quasi-isometric to a lattice, as we will later discuss in Section 6. Note that the homophily scores, as reported in Table 1, are reasonably high since nodes with similar $\hat{\varepsilon}_{k}$ tend to cluster together.
@@
New challenges for learning long-range dependencies. The proposed dataset presents a significant challenge for both GNNs and GTs. In general, such long-range dependencies create a fundamental trade-off for graph learning: while increasing the number of layers in GNNs helps propagate information from distant hops, it also leads to issues such as over-smoothing (Li et al., 2018; Nt and Maehara, 2019; Rusch et al., 2023), over-squashing (Alon and Yahav, 2021; Topping et al., 2022; Di Giovanni et al., 2023), and vanishing gradients (Arroyo et al., 2025). On the other hand, GTs, which rely on attention mechanisms with quadratic computational complexity, face scalability challenges when applied to our large-scale city networks compared to learning on smaller long-ranged graphs like those in LRGB (Dwivedi et al., 2023) and other synthetic benchmarks (Zhou et al., 2025).
@@ 3 Benchmarks: From Classical GNNs to Graph Transformers
Experimental setups. We consider transductive node classification with train/validation/test splits of $10\%/10\%/80\%$ on all City-Networks, in which we evaluate various classical GNNs and GTs. Specifically, we benchmark GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCNII (Chen et al., 2020) for GNNs; and GraphGPS (Rampášek et al., 2022), Exphormer (Shirzad et al., 2023), and SGFormer (Wu et al., 2023) for GTs. In addition, an MLP is also included to reflect the importance of graph topology (or lack thereof) in our task. Lastly, we fix the data split for benchmarking purposes and report the mean and standard deviation of the test accuracy over 5 runs with different random seeds.
@@
For fair comparison purposes, we follow the latest GNN tuning technique from Luo et al. (2024) that considers residual connection, normalization, dropout, etc. Readers are also referred to Appendix D.2 for full details on our training and experimental setups.
@@
Table 2: Best baseline results ($L=16$) on City-Networks.
Baseline Type Paris Shanghai Los Angeles London
MPNNs MLP $25.5\pm 0.4$ $28.4\pm 0.6$ $24.1\pm 0.5$ $27.9\pm 0.1$
ChebNet $54.1\pm 0.2$ $66.5\pm 0.1$ $61.4\pm 0.4$ $54.7\pm 0.2$
GCN $53.2\pm 0.3$ $62.1\pm 0.2$ $58.3\pm 0.3$ $50.1\pm 0.7$
GraphSAGE $54.6\pm 0.2$ $68.3\pm 0.5$ $61.4\pm 0.3$ $55.4\pm 0.2$
GAT $51.1\pm 0.3$ $68.0\pm 0.5$ $59.5\pm 0.3$ $52.0\pm 0.3$
GCNII $51.3\pm 0.2$ $61.5\pm 0.4$ $56.0\pm 0.3$ $48.2\pm 0.3$
DropEdge $48.2\pm 0.2$ $60.8\pm 0.4$ $55.5\pm 0.3$ $45.0\pm 0.3$
GTs GraphGPS $52.1\pm 0.6$ $63.0\pm 0.5$ $59.8\pm 0.5$ OOM
Exphormer $55.1\pm 0.8$ $70.2\pm 0.4$ $63.8\pm 0.6$ $49.5\pm 0.4$
SGFormer $52.0\pm 0.8$ $64.1\pm 0.3$ $60.1\pm 0.7$ $48.3\pm 0.3$
For completeness, we summarize the best baseline performance on City-Networks in Table 2, and further conducte an ablation study in Appendix D.3 that (i) tests baselines under different hidden_size; and (ii) fixes the model size at $L=16$, refrains the model from seeing beyond hop-$H$ at each node, then tests baselines at different $H$. The results are all consistent with our above findings, and the potential limitation of our experiments is discussed in Appendix F.
@@ 4 A Long-range Measurement Based on Jacobians
Given equation 2, we further define the average size $R$ of the influence-weighted receptive field as:
@@
where $H$ is the maximum number of hops to be considered. Intuitively, one can understand $R$ as measuring how far away the average unit of influence is, since the influences from distant nodes to the target node will be proportionally stronger in long-range tasks compared to those in short-range tasks. Note that the proposed $R$ bears similarity to the recent work of Bamberger et al. (2025): the definition of $R$ in equation 3 is based on the shortest path distance, which corresponds to a specific instance of the measure in  Bamberger et al. (2025). However, unlike that work which focuses on an aggregated measure of the range, our work has a particular focus on using the per-hop influence $T_{h}(v)$ as a diagnostic tool of the dependency decay, which leads to the key analysis in Figure 4 below.
Finally, we provide an analysis of the computation cost of our measurement in Appendix E and discuss its potential limitations on large dense graphs (a common challenge for Jacobian-based methods (Xu et al., 2018; Gasteiger et al., 2022) in the literature) in Appendix F.
@@
To better compare $\bar{T}_{h}$ across models and datasets, we normalize it by $\bar{T}_{0}$ and report $\bar{T}_{h}/\bar{T}_{0}$ in Figure 4. For all models, we generally observe a rapid influence drop at distant hops on Cora and ogbn-arxiv, whereas the decay is noticeably much slower on City-Networks. The patterns on Amazon-Ratings and PascalVOC-SP generally fall between the above two cases with influence concentrated on the first few hops, which corroborate our findings in Section 3. While $R$ and $\bar{T}_{h}$ are both model-dependent, the above results yield a consistent ordering across datasets: models trained on City-Networks are required to leverage information from more distant hops to perform well, compared to those trained on existing datasets.
Note that the bias in the models will naturally lead to a biased estimation of the ground-truth range using the proposed measurement, and we will discuss this limitation in more detail in Appendix F. In addition, we further provide results on RingTransfer (Bodnar et al., 2021), a synthetic experiment for testing long-range dependency using small ring-like graphs under inductive settings in Appendix D.4; and provide results on other common heterophilic datasets in Appendix D.5.
For the rest of the paper, we will first provide a theoretical justification of the proposed dataset by linking over-smoothing to algebraic connectivity and network diameter in Section 5, and then present the intuition behind the proposed long-range measurement in Section 6.